OpenStudy (anonymous):

We were given this statement about Rates of Change and told it is False but I don't understand why. "If we blow up a balloon so that the volume changes with a constant rate, then the radius (r(t)) changes with a constant rate." Thanks :)

OpenStudy (anonymous):

v= 4/3 pi r^3 *derivative* dv/dt = 4 pi r^2 dr/dt

OpenStudy (anonymous):

but we weren't given the equation for V so how is it false for every function?

OpenStudy (anonymous):

I'm assuming it's a sphere because I don't really know any other 3D shapes that are usually described with a radius

OpenStudy (anonymous):

Im still a bit confused...sorry...how does taking the deriv. of V with respect to t show us that? Sorry..

OpenStudy (anonymous):

I have...but I just can't see what your saying. to find the ROC of the volume wouldn't you do deriv wrt r. Why t?

OpenStudy (anonymous):

oh :)

OpenStudy (anonymous):

So are you saying... for V' to be a constant (V'=4pir^2) then r(t) needs to be a constant. and if r(t) is a constant then r'(t) will be = 0

OpenStudy (anonymous):

dv/dt = constant; that is one of the givens dv/dt = 4 pi r^2 dr/dt dv/dt being constant doesn't NECESSARILY imply that dr/dt is constant. You can find counterexamples for which a having a constant dv/dt doesn't entail any specific dr/dt because there is also a r^2 term

OpenStudy (anonymous):

dv/dt = 4 pi r^2 * dr/dt (var1) = 4 pi *(var2)^2 * (var3) there are three variables in the equation so even if you know one variable (dv/dt), you can't say one of the other variables (dr/dt) is constant because the equation is also dependent on a third variable (r)

OpenStudy (anonymous):

oh ok thanks :)